# Domain-Invariant Representation Learning of Bird Sounds

### v6

@@ 1 Introduction
Despite its effectiveness, SupCon is computationally expensive due to the need for pairwise similarity calculations between all examples. To address this, in this paper we introduce ProtoCLR (Prototypical Contrastive Learning of Representations), a more efficient variant. By analyzing SupCon’s gradient and drawing inspiration from the generalization capabilities of prototypical learning [11, 12] in few-shot learning, ProtoCLR replaces pairwise comparisons with a prototypical contrastive loss that compares examples to class prototypes (one prototype per class), retaining the original objective while significantly reducing computational complexity.
We propose a new few-shot classification evaluation based on the BIRB [13] benchmark to evaluate the generalization capabilities of models trained on XC’s focal recordings and tested on diverse soundscape datasets. This benchmark is designed to assess how well models can generalize across domains in challenging few-shot scenarios. We validate our proposed loss ProtoCLR on this benchmark, demonstrating its effectiveness in improving DG in bird sound classification.
@@ 3.1 Supervised Contrastive Loss (SupCon)
This gradient consists of two terms: a positive term $\frac{1}{|P(i)|}\sum_{p\in P(i)}\frac{\displaystyle 1}{\displaystyle\tau}{z}_{p}$ pulling the embedding ${z}_{i}$ towards its class centroid and a negative term $-\frac{\displaystyle 1}{\displaystyle\tau}\frac{\sum_{a\in A(i)}S_{ia}{z}_{a}}{\sum_{a\in A(i)}S_{ia}}$ pushing it away from other examples.
@@ 4 Experiments
BirdAVES-biox-base 95 7.41$\pm$1.0 26.4$\pm$2.3 13.2$\pm$3.1 9.84$\pm$3.5 8.74$\pm$0.6 14.1$\pm$3.1 13.2
BirdAVES-bioxn-large 316 7.59$\pm$0.8 27.2$\pm$3.6 13.7$\pm$2.9 12.5$\pm$3.6 10.0$\pm$1.4 14.5$\pm$3.2 14.2
BioLingual 153 6.21$\pm$1.1 37.5$\pm$2.9 17.8$\pm$3.5 17.6$\pm$5.1 22.5$\pm$4.0 26.4$\pm$3.4 21.3
Perch 80 9.10$\pm$5.3 42.4$\pm$4.9 19.8$\pm$5.0 26.7$\pm$9.8 22.3$\pm$3.3 29.1$\pm$5.9 24.9
@@
BirdAVES-biox-base 95 11.6$\pm$0.8 39.7$\pm$1.8 22.5$\pm$2.4 22.1$\pm$3.3 16.1$\pm$1.7 28.3$\pm$2.3 23.3
BirdAVES-bioxn-large 316 15.0$\pm$0.9 42.6$\pm$2.7 23.7$\pm$3.8 28.4$\pm$2.4 18.3$\pm$2.3 27.3$\pm$2.3 25.8
BioLingual 153 13.6$\pm$1.3 65.2$\pm$1.4 31.0$\pm$2.9 34.3$\pm$3.5 43.9$\pm$0.9 49.9$\pm$2.3 39.6
Perch 80 21.2$\pm$1.2 71.7$\pm$1.5 39.5$\pm$3.0 52.5$\pm$5.9 48.0$\pm$1.9 59.7$\pm$1.8 48.7
@@ 4.4 Results and Discussion
Table 2 presents the performance of different models on 1-shot and 5-shot bird sound classification tasks. ProtoCLR outperforms SupCon in both 1-shot and 5-shot classification.

### v7

@@ 1 Introduction
Despite its effectiveness, SupCon is computationally expensive due to the need for pairwise similarity calculations between all examples. To address this, in this paper we introduce ProtoCLR (Prototypical Contrastive Learning of Representations), a more efficient variant. By analyzing SupCon's gradient and drawing inspiration from the generalization capabilities of prototypical learning [11, 12] in few-shot learning, ProtoCLR replaces pairwise comparisons with a prototypical contrastive loss that compares examples to class prototypes (one prototype per class), retaining the original objective while significantly reducing computational complexity.
We propose a new few-shot classification evaluation based on the BIRB [13] benchmark to evaluate the generalization capabilities of models trained on XC's focal recordings and tested on diverse soundscape datasets. This benchmark is designed to assess how well models can generalize across domains in challenging few-shot scenarios. We validate our proposed loss ProtoCLR on this benchmark, demonstrating its effectiveness in improving DG in bird sound classification.
@@ 3.1 Supervised Contrastive Loss (SupCon)
This gradient consists of two terms: a positive term $\frac{1}{|P(i)|}\sum_{p\in P(i)}\frac{\displaystyle 1}{\displaystyle\tau}{z}_{p}$ pulling the embedding ${z}_{i}$ towards its class centroid and a negative term $\frac{\displaystyle 1}{\displaystyle\tau}\frac{\sum_{a\in A(i)}S_{ia}{z}_{a}}{\sum_{a\in A(i)}S_{ia}}$ pushing it away from other examples.
@@ 4 Experiments
[…]
@@
[…]
@@ 4.4 Results and Discussion
Table 2 presents the performance of different models on 1-shot and 5-shot bird sound classification tasks. ProtoCLR outperforms SupCon in both settings, while also being more computationally efficient during training: for 1 epoch on XC with a batch size of 256, SupCon requires 80.4B multiply-accumulate operations, whereas ProtoCLR requires only 28.3B.